Graph the inequality.the sign is less than or equal to
x – y \< 2

Answer by tutorcecilia(2152)

x-y<=2
x-x-y<=-x+2
-y<=-x+2
-y/-1=>(-x/-1)+(2/-1) [flip the sign when multiplying or dividing by a negative number]
.
y=>x-2
To graph, use a solid line for the boundary line; shade everything above the line.
Graphing: